there are five people standing in a room. Antonio is 5 time the age of Patrick who is half the age of Merlyn. Ed is 30 years younger than double Merlyn and Patrick's combined ages. Precious is 79 years younger than Antonio. The sum of their ages is 271. How old is Precious?
Answer by mananth(16946)

there are five people standing in a room. Antonio is 5 time the age of Patrick who is half the age of Merlyn. Ed is 30 years younger than double Merlyn and Patrick's combined ages. Precious is 79 years younger than Antonio. The sum of their ages is 271. How old is Precious?
..
let Antonio = x years
Precious = x-79
Antonio = x= 5* age of patrick
Patrick = x/5
..
Patrick = 1/2 age of Merlyn
2*patrick = Merlyn
2x/5= Merlyn
..
Ed = 2*(2x/5 + x/5)-30
Ed = 2*(3x/5)-30
Ed= 6x/5-30
...
now you have all their ages in variables
add them up
x+x-79+x/5 +2x/5 +6x/5-30 =271
2x+x/5 +2x/5+6x/5 = 271+79+30
(10x+x+2x+6x)/5 =380
19x=5*380
x= 5*380/19
x = 100 Antonio's age
Precious = Antonio's age -79
..
100-79 = 21
Precious is 21 years
..
m.ananth@hotmail.ca